### v1

@@ 5.3 Empirical Comparison of Optimization Algorithms
In the final experiment, we revisit the regularized low-rank problem from Fig. 3 to compare AML–Newton (random) with the RSN method [11, 33]. Prior work shows that subspace and multilevel methods are well suited to this problem class and can outperform state-of-the-art methods [17, 36]. With this comparison, our aim is to demonstrate that AML–Newton further improves the performance of second-order methods. Like AML–Newton (random), RSN will generate the random subspaces according to Definition 5.1 at each iteration. For AML–Newton we set $m=2$ models, that is one coarse and one fine. This configuration makes AML–Newton closely related to Randomized Subspace Newton, with the key difference that it dynamically chooses whether to compute directions in the subspace or in the full space. The result regarding the convergence behavior of the methods are presented in Fig. 6, and Table 6 shows the number of coarse and fine steps computed by AML–Newton during the optimization.

### v2

@@ 5.3 Empirical Comparison of Optimization Algorithms
In the final experiment, we revisit the regularized low-rank problem from Fig. 3 to compare AML–Newton (random) with the RSN method [gower2019rsn, tsipinakis2024multilevel]. Prior work shows that subspace and multilevel methods are well suited to this problem class and can outperform state-of-the-art methods [ho2019newton, tsipinakis2023lowrank]. With this comparison, our aim is to demonstrate that AML–Newton further improves the performance of second-order methods. Like AML–Newton (random), RSN will generate the random subspaces according to Definition 5.1 at each iteration. For AML–Newton we set $m=2$ models, that is one coarse and one fine. This configuration makes AML–Newton closely related to Randomized Subspace Newton, with the key difference that it dynamically chooses whether to compute directions in the subspace or in the full space. The result regarding the convergence behavior of the methods are presented in Fig. 6, and Table 6 shows the number of coarse and fine steps computed by AML–Newton during the optimization.